please explain the reasons and show the specific calculation process
On January 1, 2009, ABC Inc issued $1,000,000 of bonds: -The bonds are due on December 31, 2018. -Interest on the bonds is payable on June 30th and December 31st. The annual effective rate for the debt was 8% (4% per six months) -The annual coupon rate for the debt was 12% (6% per six months) - The company receive -Legal and other costs of $12,000 were incurred in connection with the issue. -The company uses the effective interest method to amortize any discount or premium -On January 1, 2010 the company retires 20% of the bonds paying $260,000. d S1,271807 from bond investors. What entry does ABC record on January 1,2010?
Answer #1

Answer:

Issued on January 1, 2009:

Face value of bond =$1,000,000

Interest payable on June 30 and Dec 31

Effective interest rate for six months = 4%

Coupon rate for six months = 6%

Bond sale value = $1,271,807

Bond Premium = $1,271,807 - $1,000,000 =$271,807

Uses effective interest method to amortize premium or discount

June 30, 2009:

Interest expense = $1,271,807 * 4% = $50,872.28

Cash paid to bond investors = $1,000,000 * 6% = $60,000

Amortization premium = $60,000 - $50,872.28 = 9127.72

Carrying value of bond = $1,271,807 - 9127.72 = $1,262,679.28

December 31, 2009:

Interest expense = $1,262,679.28 * 4% = $50,507.17

Cash paid to bond investors = $1,000,000 * 6% = $60,000

Amortization premium = $60,000 - $50,507.17 = $9,492.83

Carrying value of bond:

Face value = $1,000,000

Premium = $271,807 - 9127.72 - $9,492.83 = $253,186.45

Carrying value = $1,000,000 + $253,186.45 = $1,253,186.45

January 1, 2010:

20% of bonds retired paying $260,000

Carrying value of 20% of bonds = $1,253,186.45 * 20% = $250,637.29

Face value of bond retired = 1,000,000 * 20% = $200,000

Un-amortized premium of bonds retired = $253,186.45 * 20% = $50,637.29

Loss on retirement of bonds = $260,000 - $250,637.29 = $9,362.71

Journal entry to record on January 1, 2010 is:

DebitCredit Account Title and explanation Bonds payable Premium on bonds Loss on retirement of bonds Date $200,000.00 $50,637